# RLHS: Mitigating Misalignment in RLHF with Hindsight Simulation

### v1

@@ Abstract
Generative AI systems like foundation models (FMs) must align well with human values to ensure their behavior is helpful and trustworthy. While Reinforcement Learning from Human Feedback (RLHF) has shown promise for optimizing model performance using human judgments, existing RLHF pipelines predominantly rely on immediate feedback, which can fail to accurately reflect the downstream impact of an interaction on users’ utility. We demonstrate that feedback based on evaluators’ foresight estimates of downstream consequences systematically induces Goodhart’s Law dynamics, incentivizing misaligned behaviors like sycophancy and deception and ultimately degrading user outcomes. To alleviate this, we propose decoupling evaluation from prediction by refocusing RLHF on hindsight feedback. Our theoretical analysis reveals that conditioning evaluator feedback on downstream observations mitigates misalignment and improves expected human utility, even when these observations are simulated by the AI system itself. To leverage this insight in a practical alignment algorithm, we introduce Reinforcement Learning from Hindsight Simulation (RLHS), which first simulates plausible consequences and then elicits feedback to assess what behaviors were genuinely beneficial in hindsight. We apply RLHS to two widely-employed online and offline preference optimization methods—Proximal Policy Optimization (PPO) and Direct Preference Optimization (DPO)—and show empirically that misalignment is significantly reduced with both methods. Through an online human user study, we show that RLHS consistently outperforms RLHF in helping users achieve their goals and earns higher satisfaction ratings, despite being trained solely with simulated hindsight feedback. These results underscore the importance of focusing on long-term consequences, even simulated ones, to mitigate misalignment in RLHF.
@@ 1 Introduction
Aligning artificial intelligence (AI) systems with human values and intentions is crucial to ensuring they behave in ways that are helpful, honest, and trustworthy. A widely-deployed method for achieving this alignment is through human feedback (Leike et al., 2018), with successful applications to, e.g., training AI assistants (Glaese et al., 2022; Touvron et al., 2023; Anthropic, 2023; Achiam et al., 2023). In particular, Reinforcement Learning from Human Feedback (RLHF) (Christiano et al., 2017; Ziegler et al., 2019; Ouyang et al., 2022; Stiennon et al., 2020) leverages human feedback to fine-tune and align foundation models (FMs). While RLHF has shown promise in aligning models with human preferences, it often relies heavily on human perceptions during interactions, which may not accurately reflect the downstream consequences of the service provided. Such myopic feedback can misguide the model’s behavior and limit its effectiveness in aligning with human values. For example, human evaluators could misjudge an interaction on the spot, due to limited resources (e.g., partial observability; Casper et al. 2023; Lang et al. 2024) or limited bandwidth (e.g., constraints on time, attention, or care; Pandey et al. 2022; Chmielewski & Kucker 2020), leading to incomplete or misinformed feedback. A recent study has theorized that partial observability of an AI assistant’s task execution during human–AI interaction can lead RLHF to learn deceptive behaviors (Lang et al., 2024).
Figure 1: RLHF can incentivize AI systems to provide inaccurate or deceptive information to their users, prioritizing positive on-the-spot feedback and neglecting long-term consequences. For example, a customer may prefer to hear good news while shopping but will ultimately be disappointed (and objectively worse off) if stuck with an ill-informed purchase. The proposed RLHS introduces hindsight for human feedback, focusing on evaluations after simulating the outcome. This enables more informed feedback that improves alignment between the AI and the human’s true utility.
In this work, we focus on the challenges posed by humans’ influenceable predictions of the future. In many settings, the utility provided by an AI system to a human user (and similarly its “helpfulness” and “harmlessness”, which RLHF evaluators are typically asked to assess) is not an intrinsic property of the outputs that it generates but rather a function of their real-world consequences, brought about by the user’s real-world decisions upon consuming said outputs. We hypothesize that relying on human users’ prediction of the helpfulness of an interaction right after it takes place creates a pernicious Goodhart’s law dynamic: incentivizing the AI system to increase users’ subjective foresight value will favor inducing unrealistically optimistic expectations in users—while at best these may be innocuous, at worst they can lead users to make poor choices resulting in degraded outcomes.
@@
Our key finding is that equipping evaluator feedback with the benefit of hindsight—even if this is simulated using imperfect models—can significantly reduce model misalignment with the evaluator’s true utility, decreasing the chances of deceptive and misleading outputs. We implement hindsight simulation with both offline and online preference optimization approaches, including direct preference optimization (DPO) (Rafailov et al., 2024) and proximal policy optimization (PPO) (Schulman et al., 2017) and show empirically that it greatly improves alignment in both training paradigms. We also present results from human user studies, in which RLHS consistently improves both users’ ground-truth utility and subjective satisfaction, despite being trained with only simulated hindsight feedback. Our comparative findings demonstrate that RLHS significantly outperforms non-hindsight methods—specifically Reinforcement Learning from AI Feedback (RLAIF), which similarly uses AI generation as a proxy for real human feedback, and has been shown to produce results closely resembling that of RLHF (Bai et al., 2022b; Lee et al., 2023).
@@ 2 Background and Preliminaries
Human Decision-Making under Uncertainty. We consider a decision problem faced by a human entity (e.g., an individual, group, or institution) under predictive uncertainty and imperfect observations. We can model such a problem as a partially observable Markov decision process (POMDP) defined by a tuple $\mathcal{P}^{\textit{H}}=(\mathcal{S},\mathcal{A}^{\textit{H}},\mathcal{O}^{\textit{H}},\mathcal{T},{O}^{\textit{H}},P_{0},r,\gamma,{\theta}^{\textit{H}})$, where $\mathcal{S}$ is the set of relevant world states, $\mathcal{A}^{\textit{H}}$ is the set of available actions, $\mathcal{O}^{\textit{H}}$ is the human’s observation space, $\mathcal{T}:\mathcal{S}\times\mathcal{A}^{\textit{H}}\rightarrow\Delta(\mathcal{S})$ is the stochastic transition kernel, ${O}^{\textit{H}}:\mathcal{S}\rightarrow\Delta(\mathcal{O}^{\textit{H}})$ is the human’s observation map, $P_{0}\in\Delta(\mathcal{S})$ is the initial state distribution, $r:\mathcal{S}\times\mathcal{A}^{\textit{H}}\times\Theta^{\textit{H}}\rightarrow\mathbb{R}$ is the reward function, $\gamma\in(0,1)$ is the time discount factor, and ${\theta}^{\textit{H}}\in\Theta^{\textit{H}}$ describes the human’s intrinsic preferences. Due to partial observability of the world state ${s}\in\mathcal{S}$, the human may maintain an internal state ${z}^{\textit{H}}\in{\mathcal{Z}^{\textit{H}}}$ (e.g., a belief $b^{H}\in\Delta(\mathcal{S})$ encoding the human’s uncertain knowledge of the world state, although ${z}^{\textit{H}}$ may be thought of as a more general variable that could encode features such as the human’s emotional state or attention focus). The human may be modeled as taking actions according to a stochastic policy ${\pi}^{\textit{H}}:{\mathcal{Z}^{\textit{H}}}\rightarrow\Delta(\mathcal{A}^{\textit{H}})$.
@@
AI-Assisted Human Decision-Making. When the human consults an AI system (e.g., a FM) to help with their decision problem, we may augment the above problem with the human–AI interaction. The resulting Assisted POMDP is a tuple $\mathcal{P}^{\textit{H}}_{\rightleftharpoons}=(\mathcal{S},\mathcal{A}^{\textit{H}}\times\mathcal{A}^{\textit{H}}_{\rightleftharpoons},\mathcal{A}^{\textit{AI}}_{\rightleftharpoons},\mathcal{O}^{\textit{H}},\mathcal{O}^{\textit{AI}},\mathcal{T},{O}^{\textit{H}},{O}^{\textit{AI}},P_{0},r,\gamma,{\theta}^{\textit{H}})$, where $\mathcal{A}^{\textit{H}}_{\rightleftharpoons}$ and $\mathcal{A}^{\textit{AI}}_{\rightleftharpoons}$ are the sets of interactive actions available to the human and AI system, $\mathcal{O}^{\textit{AI}}$ is the AI’s observation space, and ${O}^{\textit{AI}}$ is the AI’s observation map ${O}^{\textit{AI}}:\mathcal{S}\rightarrow\Delta(\mathcal{O}^{\textit{AI}})$. In this model, the AI takes an advisory role: it can respond to a human’s interactive action ${a}_{\rightleftharpoons}^{\textit{H}}\in\mathcal{A}^{\textit{H}}_{\rightleftharpoons}$ (e.g., a query through a chat interface) with its own ${a}^{\textit{AI}}_{\rightleftharpoons}\in\mathcal{A}^{\textit{AI}}_{\rightleftharpoons}$ (e.g., a generated text or multimedia output). After one or multiple rounds of such interactions, the human may take a physical action ${a}^{\textit{H}}\in\mathcal{A}^{\textit{H}}$ to affect the evolution of the world state ${s}$. This Assisted POMDP is a special case of a partially observable stochastic game (POSG) (Hansen et al., 2004). In such interactions, the AI’s goal is to influence the human’s internal state ${z}^{\textit{H}}$ towards maximizing the rewards $r({s},{a}^{\textit{H}};\theta)$ accrued over time. This, however, is made challenging by the AI’s fundamental uncertainty about the human’s preferences ${\theta}^{\textit{H}}$.
@@
Reinforcement Learning from Human Feedback (RLHF). RLHF aims to learn the human’s preferences ${\theta}^{\textit{H}}$ from human feedback data, which typically involves three key steps. In Step 1, the human is asked to provide feedback on some state sequences ${\mathbf{{s}}}=({s}_{0},{s}_{1},\ldots,{s}_{T})$ (e.g., a human–AI dialogue), with ${s}_{t}\in\mathcal{S},~{}\forall t=0,1,\ldots,T$. For example, in binary comparison (Christiano et al., 2017), assuming human is a Boltzmann rational decision maker (Luce, 1959), the probability that the human prefers ${\mathbf{{s}}}$ over ${\mathbf{{s}}}^{\prime}$ is $P^{r}_{T}({\mathbf{{s}}}\succ{\mathbf{{s}}}^{\prime})=\sigma(\beta(R_{T}({\mathbf{{s}}})-R_{T}({\mathbf{{s}}}^{\prime})))$, where $\sigma(\cdot)$ is the sigmoid function, $\beta>0$ is the inverse temperature parameter, and $R_{T}({\mathbf{{s}}})=\sum_{t=0}^{T}\gamma^{t}r({s}_{t})$ is the return received by state sequence ${\mathbf{{s}}}$.. Step 2 is to fit a reward function $\hat{r}$ based on a dataset containing state sequences paired with human feedback, aiming for $\hat{r}$ to resemble $r$ as closely as possible. Step 3 is to compute an AI policy $\hat{{\pi}}:\mathcal{S}\rightarrow\Delta(\mathcal{A}^{\textit{AI}}_{\rightleftharpoons})$ that maximizes the return based on the estimated reward $\hat{r}$, i.e., $\hat{{\pi}}=\operatorname*{arg\,max}_{\pi}{U}_{T}({\pi})$, where ${U}_{T}({\pi}):=\operatorname*{\mathbb{E}}_{{\mathbf{{s}}}\sim p^{\pi}}[\hat{R}_{T}({\mathbf{{s}}})]$ is the expected utility of ${\pi}$, and $p^{\pi}$ is the on-policy distribution of state sequence ${\mathbf{{s}}}$ under $P_{0}$, $\mathcal{T}$, and ${\pi}$. Due to the lack of an analytical model for $\mathcal{T}$ and the high-dimensional nature of aligning modern AI models, reinforcement learning (RL) is often used to approximately optimize the policy at scale. Recent studies have revealed that RLHF can lead to misalignment when the human gives feedback based on partial observations ${\mathbf{{o}}}^{\textit{H}}=({o}^{\textit{H}}_{0},{o}^{\textit{H}}_{1},\ldots,{o}^{\textit{H}}_{T})$ rather than the previously assumed—but rarely realistic—full state sequences, resulting in deceptive or manipulative AI behaviors (Casper et al., 2023; Lang et al., 2024). We argue that RLHF misalignment more generally emerges in settings with significant human uncertainty, whether perceptual, predictive, or a combination of the two. We propose to take advantage of the general insight that assessments about past outcomes that the evaluator has experienced would be significantly less uncertain (and thus less influenceable) than assessments about future outcomes that are yet to unfold.
3 Alignment Algorithm: RL from Hindsight Simulation
To address the misalignment caused by human uncertainty in RLHF, we introduce Reinforcement Learning from Hindsight Simulation (RLHS). Our central contention is that by decoupling human feedback on the downstream outcomes of an interaction from the prediction of these outcomes, the learned human reward model and corresponding AI policy will be substantially better aligned.
@@ 3.1 Hindsight Mitigates Misalignment
Given a predictive model of the human, the AI’s decision problem in the Assisted POMDP game $\mathcal{P}^{\textit{H}}_{\rightleftharpoons}$ in Section 2 can be reformulated as a POMDP $\mathcal{P}^{\textit{AI}}_{\rightleftharpoons}=(\bar{\mathcal{S}},\mathcal{A}^{\textit{AI}}_{\rightleftharpoons},\bar{\mathcal{O}}^{\textit{AI}},\bar{\mathcal{T}},\bar{O}^{\textit{AI}},\bar{P}_{0},\bar{r},\gamma)$, where $\bar{\mathcal{S}}=\mathcal{S}\times\Theta^{\textit{H}}\times{\mathcal{Z}^{\textit{H}}}$, $\bar{\mathcal{O}}^{\textit{AI}}=\mathcal{O}^{\textit{AI}}\times\mathcal{A}^{\textit{H}}_{\rightleftharpoons}$, $\bar{\mathcal{T}}=(\mathcal{T},\mathcal{T}_{\theta},\mathcal{T}^{\textit{H}})$, $\bar{P}_{0}\in\Delta(\bar{\mathcal{S}})$, and $\bar{r}({s},{z}^{\textit{H}},{\theta}^{\textit{H}})=\operatorname*{\mathbb{E}}_{{a}^{H}\sim{\pi}^{\textit{H}}(\cdot\mid{z}^{\textit{H}})}r({s},{a}^{\textit{H}};{\theta}^{\textit{H}})$. Here, $\mathcal{T}^{\textit{H}}:{\mathcal{Z}^{\textit{H}}}\times\mathcal{A}^{\textit{AI}}_{\rightleftharpoons}\rightarrow{\mathcal{Z}^{\textit{H}}}$ is the transition kernel of the human’s internal state, modeling how the human’s knowledge about the world state is evolved based on new observations and interactions with the AI; we treat ${\theta}^{\textit{H}}$ as a constant for the purposes of this paper, with $\mathcal{T}_{\theta}$ as the identity map. Finally, $\pi^{\textit{H}}:{\mathcal{Z}^{\textit{H}}}\to\Delta(\bar{\mathcal{A}}^{\textit{H}})$, with $\bar{\mathcal{A}}^{\textit{H}}:=\mathcal{A}^{\textit{H}}\times\mathcal{A}^{\textit{H}}_{\rightleftharpoons}$. In practice the human model can be a black box (e.g., a web-trained FM). Due to the complexity of POMDP $\mathcal{P}^{\textit{AI}}_{\rightleftharpoons}$, we aim to solve it approximately using RL with hindsight feedback provided by an evaluator, which we explain in detail below.
Since the human’s utility is inherited from their original decision problem $\mathcal{P}^{\textit{H}}$, the expected utility generated by an AI policy ${\pi}^{\textit{AI}}$ is the expected return achieved by the human’s course of action. For the purposes of RLHF, we can assume that the human begins taking physical actions after the interaction:
${U}^{\textit{H}}({\pi}^{\textit{AI}}):=\operatorname*{\mathbb{E}}_{\begin{subarray}{c}\bar{s}_{0}\sim\bar{P}_{0},\quad\,\mathcal{T}^{\textit{H}}(\cdot\mid{z}^{\textit{H}}_{\tau},~{}{a}^{\textit{AI}}_{{\rightleftharpoons},\tau}),~{}\\ {a}^{\textit{H}}_{\tau}\sim{\pi}^{\textit{H}}(\cdot\mid{z}^{\textit{H}}_{\tau}),\;{a}^{\textit{AI}}_{{\rightleftharpoons},\tau}\sim{\pi}^{\textit{AI}}(\cdot\mid z^{\textit{AI}}_{\tau})\end{subarray}}\left[\sum_{t=T+1}^{\infty}\gamma^{t}r({s}_{t},{a}^{H}_{t};{\theta}^{\textit{H}})\right],$ (1)
where $t=0,1,\ldots,T$ is the human–AI interaction phase and $t=T+1,T+2,\ldots,$is the downstream human acting phase. If the human evaluates their utility immediately after the human–AI interaction phase, relying only on predictions, this is the foresight value. If, instead, they wait until after the human acting phase and use realized outcome, this is the hindsight value. The following sections define these concepts formally.
@@
The hindsight value assessed by the human at time $k\geq 0$ is equal to the expected return received so far given the human’s available information at time $k$. In this paper we will assume that the human can accurately estimate all rewards received so far, i.e.,
$\overleftarrow{{V}}_{k}({z}^{\textit{H}}_{k})\;:=\;\mathbb{E}_{\begin{subarray}{c}{s}_{t+1}\sim\mathcal{T}(\cdot|{s}_{t},{a}^{\textit{H}}_{t}),~{}{a}^{\textit{H}}_{t}\sim{\pi}^{\textit{H}}(\cdot\mid{z}^{\textit{H}}_{t})\end{subarray}}\!\biggl{[}\sum_{t=0}^{k}\gamma^{t}\,r({s}_{t},{a}^{\textit{H}}_{t})\biggr{]},$ (2)
where $\mathcal{T}$ is the environment transition kernel, which is unaffected by the AI’s output.
Definition 2 (Foresight Value).

### v2

@@ Abstract
While Reinforcement Learning from Human Feedback (RLHF) has shown promise in aligning generative AI, we present empirical evidence that it can also cause severe, systematic misalignment. We hypothesize that this stems from evaluator feedback depending on downstream outcome predictions (foresight) that can be influenced by the AI’s output, inducing Goodhart’s law dynamics. Conversely, our theoretical analysis shows that conditioning evaluator feedback on downstream observations (hindsight) inhibits this effect by decoupling the alignment signal from potentially compromised predictions—crucially, the result holds even if the observed outcomes are sampled from the AI’s own world model. Building on this insight, we introduce Reinforcement Learning from Hindsight Simulation (RLHS), which presents plausible simulated outcomes to evaluators before eliciting feedback. We demonstrate RLHS on online (PPO) and offline (DPO) large language model fine-tuning, obtaining superior alignment over RLHF in controlled consultancy-type experiments and user studies. We evaluate post-hoc on the TruthfulQA benchmark and find that, even after single-task fine-tuning, both RLHF misalignment and RLHS alignment carry over to substantially different settings.
Machine Learning, ICML
@@ 1 Introduction
Aligning artificial intelligence (AI) systems with human values and goals is crucial to ensuring their behavior is helpful, honest, and trustworthy. Eliciting human feedback is a widely-used alignment strategy (Leike et al., 2018), with successful applications to, e.g., training AI assistants (Glaese et al., 2022; Touvron et al., 2023; Anthropic, 2023; Achiam et al., 2023). In particular, Reinforcement Learning from Human Feedback (RLHF) (Christiano et al., 2017; Ziegler et al., 2019; Ouyang et al., 2022; Stiennon et al., 2020) leverages human feedback to fine-tune and align foundation models (FMs). While RLHF has shown promise in aligning models with human preferences, it relies predominantly on immediate assessments of isolated interactions, which may not accurately account for their downstream outcomes. Inaccurate user or evaluator feedback can misguide the model’s behavior and undermine the alignment process. For example, human evaluators could misjudge an interaction on the spot, due to limited knowledge (Casper et al., 2023; Lang et al., 2024) or cognitive resources  (Pandey et al., 2022; Chmielewski & Kucker, 2020), leading to incomplete or misinformed feedback. A recent study has theorized that partial observability of an AI assistant’s task execution during human–AI interaction can lead RLHF to learn deceptive behaviors (Lang et al., 2024).
Figure 1: RLHF can incentivize AI systems to provide inaccurate or deceptive information to their users, prioritizing positive on-the-spot feedback and neglecting long-term consequences. For example, a customer may prefer to hear good news while shopping but will ultimately be disappointed (and objectively worse off) if stuck with an ill-informed purchase. The proposed RLHS scheme removes this incentive by simulating the interaction’s downstream outcome before eliciting feedback. The resulting fine-tuned AI models show superior alignment with users’ underlying utility.
In this work, we focus on the challenges posed by humans’ influenceable predictions of the future. In many settings, the utility provided by an AI system to a human user (and similarly its “helpfulness” and “harmlessness”, which RLHF evaluators are typically asked to assess) is not an intrinsic property of the outputs that it generates but rather a function of their real-world consequences, brought about by the user’s real-world decisions upon consuming said outputs. We contend that rewarding an AI system to improve users or evaluators’ in-the-moment assessments of interactions creates a pernicious Goodhart’s law dynamic: such an incentive will favor outputs that induce unrealistically optimistic expectations in users—while at best these may be innocuous, at worst they can lead users to make poor choices resulting in degraded outcomes.
@@
Our key finding is that equipping evaluators with the benefit of (simulated) hindsight—and thereby relieving them of the burden of prediction—significantly reduces model misalignment after fine-tuning. Intuitively, even if the world model generating outcomes contains inaccuracies, these are not dependent on the AI output, and therefore the AI agent has no incentive to manipulate them. We evaluate hindsight simulation with both offline and online preference optimization approaches, including direct preference optimization (DPO) (Rafailov et al., 2024) and proximal policy optimization (PPO) (Schulman et al., 2017) and find that it greatly improves alignment in both training paradigms. We validate these results through an online user study, in which RLHS consistently improved the objective utility and subjective satisfaction of our human participants, despite being trained with only simulated hindsight feedback. Our comparative findings demonstrate that RLHS significantly outperforms non-hindsight methods—specifically Reinforcement Learning from AI Feedback (RLAIF), which similarly uses AI generation as a proxy for human feedback, and has been shown to produce results similar to RLHF (Bai et al., 2022b; Lee et al., 2023). Finally, a post-hoc evaluation on the TruthfulQA benchmark indicates that, even after single-task fine-tuning, both RLHF misalignment and RLHS alignment can persist in substantially different settings.
@@ 2 Background and Preliminaries
Human Decision-Making under Uncertainty. We consider a decision problem faced by a human entity (e.g., an individual, group, or institution) under predictive uncertainty and imperfect observations. We model the a problem as a partially observable Markov decision process (POMDP) defined by a tuple $\mathcal{P}^{\textit{H}}=(\mathcal{S},\mathcal{A}^{\textit{H}},\mathcal{O}^{\textit{H}},\mathcal{T},{O}^{\textit{H}},P_{0},r,\gamma,{\theta}^{\textit{H}})$, where $\mathcal{S}$ is the set of relevant world states, $\mathcal{A}^{\textit{H}}$ is the set of available actions, $\mathcal{O}^{\textit{H}}$ is the human’s observation space, $\mathcal{T}:\mathcal{S}\times\mathcal{A}^{\textit{H}}\rightarrow\Delta(\mathcal{S})$ is the stochastic transition kernel, ${O}^{\textit{H}}:\mathcal{S}\rightarrow\Delta(\mathcal{O}^{\textit{H}})$ is the human’s observation map, $P_{0}\in\Delta(\mathcal{S})$ is the initial state distribution, $r:\mathcal{S}\times\mathcal{A}^{\textit{H}}\times\Theta^{\textit{H}}\rightarrow\mathbb{R}$ is the reward function, $\gamma\in(0,1)$ is the time discount factor, and ${\theta}^{\textit{H}}\in\Theta^{\textit{H}}$ describes the human’s intrinsic preferences. Due to partial observability of the world state ${s}\in\mathcal{S}$, the human may maintain an internal state ${z}^{\textit{H}}\in{\mathcal{Z}^{\textit{H}}}$ (e.g., a belief $b^{H}\in\Delta(\mathcal{S})$) encoding the human’s uncertain knowledge of the world state, although ${z}^{\textit{H}}$ may be thought of as a more general variable that could encode features such as the human’s emotional state or attention focus). The human may be modeled as taking actions according to a stochastic policy ${\pi}^{\textit{H}}:{\mathcal{Z}^{\textit{H}}}\rightarrow\Delta(\mathcal{A}^{\textit{H}})$.
@@
AI-Assisted Human Decision-Making. When the human consults an AI system (e.g., a FM) to help with their decision problem, we may augment the above problem with the human–AI interaction. The resulting Assisted POMDP is a tuple $\mathcal{P}^{\textit{H}}_{\rightleftharpoons}=(\mathcal{S},\mathcal{A}^{\textit{H}}\times\mathcal{A}^{\textit{H}}_{\rightleftharpoons},\mathcal{A}^{\textit{AI}}_{\rightleftharpoons},\mathcal{O}^{\textit{H}},\mathcal{O}^{\textit{AI}},\mathcal{T},{O}^{\textit{H}},{O}^{\textit{AI}},P_{0},r,\gamma,{\theta}^{\textit{H}})$, where $\mathcal{A}^{\textit{H}}_{\rightleftharpoons}$ and $\mathcal{A}^{\textit{AI}}_{\rightleftharpoons}$ are the sets of interactive actions available to the human and AI system, $\mathcal{O}^{\textit{AI}}$ is the AI’s observation space, and ${O}^{\textit{AI}}$ is the AI’s observation map ${O}^{\textit{AI}}:\mathcal{S}\rightarrow\Delta(\mathcal{O}^{\textit{AI}})$. In this model, the AI takes an advisory role: it can respond to a human’s interactive action ${a}_{\rightleftharpoons}^{\textit{H}}\in\mathcal{A}^{\textit{H}}_{\rightleftharpoons}$ (e.g., a query through a chat interface) with its own ${a}^{\textit{AI}}_{\rightleftharpoons}\in\mathcal{A}^{\textit{AI}}_{\rightleftharpoons}$ (e.g., a generated text or multimedia output). After one or multiple rounds of such interactions, the human may take a physical action ${a}^{\textit{H}}\in\mathcal{A}^{\textit{H}}$ to affect the evolution of the world state ${s}$. This Assisted POMDP is a special case of a partially observable stochastic game (POSG) (Hansen et al., 2004). In such interactions, the AI’s goal is to influence the human’s internal state ${z}^{\textit{H}}$ towards maximizing the rewards $r({s},{a}^{\textit{H}};{\theta}^{\textit{H}})$ accrued over time. This, however, is made challenging by the AI’s fundamental uncertainty about the human’s preferences ${\theta}^{\textit{H}}$.
@@
Reinforcement Learning from Human Feedback (RLHF). RLHF aims to learn the human’s preferences ${\theta}^{\textit{H}}$ from human feedback data, which typically involves three key steps. In Step 1, the human is asked to provide feedback on some state sequences ${\mathbf{{s}}}=({s}_{0},{s}_{1},\ldots,{s}_{k})$ (e.g., a human–AI dialogue), with ${s}_{t}\in\mathcal{S},~{}\forall t=0,1,\ldots,k$. For example, in binary comparison (Christiano et al., 2017), assuming human is a Boltzmann-rational decision maker (Luce, 1959), the probability that the human prefers ${\mathbf{{s}}}$ over ${\mathbf{{s}}}^{\prime}$ is $P^{r}_{k}({\mathbf{{s}}}\succ{\mathbf{{s}}}^{\prime})=\sigma(\beta(R_{k}({\mathbf{{s}}})-R_{k}({\mathbf{{s}}}^{\prime})))$, where $\sigma(\cdot)$ is the sigmoid function, $\beta>0$ is the inverse temperature parameter, and $R_{k}({\mathbf{{s}}})=\sum_{t=0}^{T}\gamma^{t}r({s}_{t})$ is the return received by state sequence ${\mathbf{{s}}}$. Step 2 is to fit a reward function $\hat{r}$ based on a dataset containing state sequences paired with human feedback, hoping that $\hat{r}$ will resemble $r$ as closely as possible. Step 3 is to compute an AI policy $\hat{{\pi}}:\mathcal{S}\rightarrow\Delta(\mathcal{A}^{\textit{AI}}_{\rightleftharpoons})$ that maximizes the return based on the estimated reward $\hat{r}$, i.e., $\hat{{\pi}}=\operatorname*{arg\,max}_{\pi}{U}_{k}({\pi})$, where ${U}_{k}({\pi}):=\operatorname*{\mathbb{E}}_{{\mathbf{{s}}}\sim p^{\pi}}[\hat{R}_{k}({\mathbf{{s}}})]$ is the expected utility of ${\pi}$, and $p^{\pi}$ is the on-policy distribution of state sequence ${\mathbf{{s}}}$ under $P_{0}$, $\mathcal{T}$, and ${\pi}$. Due to the lack of an analytical model for $\mathcal{T}$ and the high-dimensional nature of aligning modern AI models, reinforcement learning (RL) is often used to approximately optimize the policy at scale. Recent studies have revealed that RLHF can lead to deceptive AI behaviors when the human gives feedback based on partial observations  (Casper et al., 2023; Lang et al., 2024). We argue that RLHF misalignment more generally emerges in settings with significant human uncertainty, whether perceptual, predictive, or a combination of the two. We propose to take advantage of the general insight that assessments about past outcomes that the evaluator has experienced would be significantly less uncertain (and thus less influenceable) than assessments about future outcomes that are yet to unfold.
3 Algorithm: RL from Hindsight Simulation
We begin by examining how RLHF can become misaligned due to the limitations of human foresight. We then discuss the advantages of incorporating hindsight and explain why and how it mitigates misalignment. Finally, we introduce Reinforcement Learning from Hindsight Simulation (RLHS) —a principled alignment method that employs simulated hindsight to decouple feedback from potentially AI-compromised predictions.
@@ 3.1 Hindsight Mitigates Misalignment
We consider the setting in which the human interacts with an AI up to time $k$ (interaction phase), and then proceeds to take actions (acting phase). Once the human begins to take physical actions, we are interested in the human utility, defined as a discounted sum or rewards (the reward function being unknown to the AI system). We can define the utility of any given trajectory, or equivalently, its hindsight value from the end of the horizon.
@@
Let $\tau_{k:k+N}=({s}_{t},{o}^{\textit{H}}_{t},{a}^{\textit{H}}_{t})_{t=k}^{k+N}$ denote the trajectory of states, observations, and human actions from time $k\geq 0$ to $k+N$. The hindsight value of any such trajectory is
${U}(\tau_{k:k+N};{\theta}^{\textit{H}}):=\sum_{t=k}^{k+N}\gamma^{t}r({s}_{t},{a}^{\textit{H}}_{t};{\theta}^{\textit{H}})\,.$ (1)
Definition 2 (Human rollout distribution).
Let $P(\tau_{k:k+N}\mid s_{k},z^{H}_{k})$ be the probability distribution over the trajectory $\tau_{k:k+N}=(s_{t},o_{t}^{H},a_{t}^{H})_{t=k}^{k+N}$ induced by:
${a}_{t}^{\textit{H}}$ $\sim\pi^{\textit{H}}(\cdot\mid{z}^{\textit{H}}_{t}),$ $\quad{s}_{t+1}$ $\sim\mathcal{T}_{s}(\cdot\mid{s}_{t},{a}_{t}^{\textit{H}}),$
${o}^{\textit{H}}_{t+1}$ $\sim\mathcal{O}^{\textit{H}}(\cdot\mid{s}_{t},{a}_{t}^{\textit{H}}),$ $\quad z_{t+1}^{H}$ $\sim\mathcal{T}_{z}(\cdot\mid{z}^{\textit{H}}_{t},{a}_{t}^{\textit{H}},{o}^{\textit{H}}_{t+1})$
where $\mathcal{T}_{s}$ and $\mathcal{O}^{{\textit{H}}}$ constitute the world model, while $\pi^{\textit{H}}$ and $\mathcal{T}_{z}$ are the human model. [^1] Since we have defined a general POMDP, it is convenient to assume that all parties’ predictive uncertainty about the world is part of their state distribution. This allows us to treat the human and the AI as sharing the same world model but possibly with different beliefs over the state. We can then express the expected utility of the human in the POMDP from any $({s},{z}^{\textit{H}})$.
$E\,{U}({s}_{k},{z}^{\textit{H}}_{k})\;=\;\mathbb{E}_{\tau_{k:k+N}\sim P(\cdot\mid s_{k},{z}^{\textit{H}}_{k})}\bigl{[}{U}(\tau_{k:k+N})\bigr{]}$ (2)
The key mathematical insight underpinning RLHS is that, for feedback purposes, the human’s internal state at the end of the interaction ${z}^{\textit{H}}_{k}$ should only affect the predicted rollout distribution through the human’s behavior ${\pi}^{\textit{H}}(\cdot|{z}^{\textit{H}}_{t})$; it should not contaminate the prediction of any given action’s outcome. By examining the below definition, we can see that this condition is readily violated by the value approximated by immediate feedback.
Definition 3 (Human’s Foresight Value).